### v2

# Multi-Microphone and Multi-Modal Emotion Recognition in Reverberant Environment

Abstract
This paper presents a Multi-modal Emotion Recognition (MER) system designed to enhance emotion recognition accuracy in challenging acoustic conditions. Our approach combines a modified and extended Hierarchical Token-semantic Audio Transformer (HTS-AT) for multi-channel audio processing with an D Convolutional Neural Networks (CNN) model for video analysis. We evaluate our proposed method on a reverberated version of the Ryerson audio-visual database of emotional speech and song (RAVDESS) dataset using synthetic and real-world Room Impulse Responsess. Our results demonstrate that integrating audio and video modalities yields superior performance compared to uni-modal approaches, especially in challenging acoustic conditions. Moreover, we show that the multimodal (audiovisual) approach that utilizes multiple microphones outperforms its single-microphone counterpart.
Index Terms— Emotion recognition, multi-modal, reverberant conditions, audio transformer
1 Introduction
Emotion Recognition (ER) is critical in human-computer interaction, with applications ranging from healthcare to customer service. Humans naturally express emotions across multiple modalities, including facial expressions, language, speech, and gestures. Accurately modeling the interactions between these modalities, which contain complementary and potentially redundant information, is essential for effective emotion recognition. Most existing studies primarily focus on uni-modal emotion recognition, concentrating on either text, speech, or video [1, 2, 3]. Although significant advancements in single-modal emotion recognition have been demonstrated, these models often fall short in complex scenarios since they do not utilize the inherently multi-modal nature of emotional expression. Moreover, research on jointly employing multi-modal and multi-microphones for ER is relatively scarce.
Previous works have made significant strides in MER. Studies such as [4, 5, 6, 7] have developed systems that simultaneously analyze visual and acoustic data. In [8], researchers presented an unsupervised MER feature learning approach incorporating audio-visual and textual information. These studies often overlooked the challenges posed by real-world single-channel acoustic conditions, particularly reverberation and noise, which can significantly impact the performance of audio-based emotion recognition.
Feature selection is vital in designing effective MER systems. For acoustic features, log-mel filterbank energies and log-mel spectrograms have been widely adopted [9, 10]. In the video domain, various deep learning architectures such as VGG16 [11], I3D [12], and FaceNet [13] have been employed, along with facial features like landmarks and action units extracted using tools like OpenFace [14]. For the text modality, Global Vectors for Word Representation (GloVe) [15] have been frequently used [14, 16, 17].
Despite these advancements, a gap remains in addressing the challenges posed by reverberant and noisy environments. Real-world acoustic conditions can significantly alter speech signals, potentially degrading the performance of audio-based emotion recognition systems. Moreover, the integration of multi-channel audio processing techniques with video analysis for emotion recognition in such challenging conditions has not been thoroughly explored.
This work addresses these limitations by proposing a MER with multi-channel audio that outperforms solutions solely based on single-channel audio. We propose a novel approach that combines two state-of-the-art architectures for audio-visual emotion recognition. The multi-channel extension of the HTS-AT architecture for the audio modality [18] and the D CNN model [19] for the video modality.
We use a reverberated version of the RAVDESS database to analyze the proposed scheme’s performance. Reverberation was added by convolving the speech utterances with real-life RIRs drawn from the Acoustic Characterisation of Environments (ACE) challenge dataset [20].
2 Problem Formulation
Denote the two modalities as . Also, denote the set of emotions as:
Let be the video signal and the anechoic audio signal, with the time index. An array of microphones captures the audio signal after propagating in the acoustic environment. The signals, as captured by the microphones, are given by:
where are the RIRs from the source to the th microphone. The feature embeddings for each modality are denoted and , respectively. This study aims to classify the utterance to one of the emotions using the available information and utilizing the relations between the feature embeddings of both modalities:
where stands for late fusion concatenation.
3 Proposed Model
Our proposed MER architecture leverages two powerful models: the modified and extended HTS-AT [18] for multi-channel audio processing and the D model [19] for video analysis. These models are uni-modal models that are combined to create a robust multi-modal system for emotion recognition in adverse acoustic conditions.
The input features of the models are the mel-spectrograms for the audio track and the raw RGB facial images for the visual track. For the audio modality, we followed the same preprocessing procedure as in [18] and used the SpecAugment Library [21] to augment the mel-spectrograms. For augmenting the video modality, we used the TorchVision Library [22].
Audio: The extended multi-channel HTS-AT model addresses the integration of multi-microphone information, employing the Swin-Transformer architecture [23],111github.com/microsoft/Swin-Transformer a variant of the Vision Transformer (ViT) [24] architecture. The architecture is also applicable to the single-microphone configurations, namely . The model’s architecture consists of four networks, each made up of Swin-Transformer blocks with varying depths. In addition, the model uses a hierarchical structure and windowed attention mechanism to efficiently process mel-spectrograms, which serve as our audio feature extractor. We use the two multi-channel variants with the modified HTS-AT module, as proposed in [18]: 1) Patch-Embed Summation - the mel-spectrogram of each channel is processed through a shared Patch-Embed layer, after which the outputs are summed across channels; and 2) Average mel-spectrograms - mel-spectrograms from multiple channels are averaged before being fed into the model. More details can be found in [18].
Video: For video feature extraction, we employ the pre-trained D model, an 18-layer ResNet-based architecture designed for action recognition. The D model decomposes the 3D convolutions into separate spatial (2D) and temporal (1D) convolutions, which allows it to effectively capture both spatial and temporal features in the video data. Fig. 1 presents the integration of the two modalities.
The feature embeddings are extracted from the extended multi-channel HTS-AT and the D models, followed by concatenation to create a unified multi-modal representation. This combined feature vector captures both audio and visual cues relevant to emotion recognition. The concatenated features are then passed through two fully connected layers for final classification. These layers learn to interpret the combined audio-visual features and to map them to emotion categories. The output of the final layer corresponds to the predicted emotion class. This integrated scheme of both modalities ensures that both multi-channel audio and visual data are effectively processed and leveraged, allowing the model to capture and utilize complementary information from both modalities, thus achieving improved ER accuracy.
4 Experimental Study
This section outlines the experimental setup and describes the comparative analysis between the proposed scheme and a baseline method.
4.1 Dataset
Our work utilized the RAVDESS dataset for emotion recognition. This dataset includes 24 actors, equally divided between male and female speakers, each delivering 60 English sentences. Hence, there are 1440 audio-video pairs representing eight different emotions (‘sad,’ ‘happy,’ ‘angry,’ ‘calm,’ ‘fearful,’ ‘surprised,’ ‘neutral,’ and ‘disgust’). All utterances are pre-transcribed. Therefore, the emotions are expressed more artificially compared to spontaneous conversation. The RAVDESS dataset is balanced across most classes, except for the neutral class, and has a relatively small number of utterances. We used an actor-split approach, dividing the data into 80% training, 10% validation, and 10% test sets, ensuring no actor appears in more than one split. As a result, model accuracy may be lower than reported in some prior works because the test set includes actors not seen during fine-tuning.
We used synthesized RIRs for fine-tuning the model for the multi-channel experiments. We employed the ‘gpuRIR‘ Python package222github.com/DavidDiazGuerra/gpuRIR to simulate reverberant multi-channel microphone signals (setting the number of microphones to ). Each clean audio sample from the RAVDESS dataset was convolved with distinct multi-channel RIRs, resulting in 1440 3-microphone audio samples. The associated video data is unaffected by reverberation.
We simulated rooms with lengths and widths uniformly distributed between 3 m and 8 m, maintaining a constant height of 2.9 m and an aspect ratio between 1 and 1.6. We randomly positioned the sound source and microphones within these simulated environments under the following constraints. The sound source was placed at a fixed height of 1.75 m, with its and coordinates randomly determined within the room, ensuring a minimum distance of 0.5 m from the room walls. Similarly, the microphones were positioned at a fixed height of 1.6 m, with their and coordinates also randomly determined within the room dimensions. The reverberation time was set at ms. The distance between the sound source and microphones was randomly selected in the range m, where to the critical distance, determined by the room’s volume and . This ensures the dominance of direct sound over reflections. Finally, we added spatially-white noise at an signal-to-noise ratio (SNR) of dB to each reverberant signal. This noise was synthesized by applying an auto-regressive filter of order 1 to white Gaussian noise, emphasizing lower frequencies.
The proposed scheme was evaluated using real-world RIRs drawn from the ACE database [20]. The ACE RIR database comprises recordings from seven different rooms with varying dimensions and reverberation levels (see Table LABEL:ACE_results). We only used a subset of the database, recorded with a mobile phone equipped with three microphones in the ner-field scenario, which is a practical choice for real-world Speech Emotion Recognition (SER) applications. We convolved all audio utterances of the test set with the ACE RIRs to generate 3-microphone signals for each utterance.
Room ( [ms]) | Video | Audio | MER | ||||||||
Clean Single-Channel | Rev Single-Channel | Avg mel | Sum PE | Clean Single-Channel | Rev Single-Channel | Avg mel | Sum PE | ||||
Lecture 1 () | 69.4 (65.5-73.3) | 42.7 (38.8-47.2) | 60 (55.5-64.4) | 61.6 (57.7-65.5) | 61.1 (57.2-65) | 70 (66.1-73.8) | 75 (71.1-78.8) | 77.2 (73.3-81.1) | 78.3 (74.4-81.6) | ||
Lecture 2 () | * | 40.5 (36.6-45) | 55 (50.5-59.4) | 60.5 (56.1-64.4) | 58.8 (54.9-62.8) | 71.6 (67.7-75.5) | 76.1 (72.7-79.4) | 76.1 (72.7-80) | 78.3 (75-81.6) | ||
Lobby () | * | 40.5 (36.6-45) | 57.2 (53.3-61.1) | 63.3 (59.4-67.22) | 69.4 (65.5-73.3) | 68.3 (64.4-72.2) | 73.3 (69.4-77.2) | 77.2 (73.3-81.1) | 77.7 (74.4-81.1) | ||
Meeting 1 () | * | 42.7 (38.8-46.6) | 57.2 (52.7-61.6) | 60 (56.1-63.8) | 61.1 (57.2-65) | 70 (66.1-73.8) | 72.7 (68.8-76.6) | 77.2 (73.8-81.1) | 78.8 (75.5-82.2) | ||
Meeting 2 () | * | 38.8 (34.4-42.7) | 56.1 (51.6-60.5) | 62.7 (58.8-66.6) | 59.4 (55.5-63.8) | 71.1 (67.2-75) | 75 (71.1-78.8) | 78.8 (75.5-82.2) | 75.5 (71.6-78.8) | ||
Office 1 () | * | 42.7 (38.3-46.6) | 59.4 (55-63.3) | 62.7 (58.8-68.3) | 63.3 (58.8-67.2) | 70 (66.1-73.8) | 75 (71.1-78.8) | 77.7 (74.4-81.6) | 77.2 (73.8-80.5) | ||
Office 2 () | * | 47.7 (43.3-52.2) | 56.6 (52.2-60.5) | 64.4 (60.5-68.3) | 60.5 (56.6-64.4) | 68.8 (65-72.7) | 70 (66.1-73.8) | 74.4 (70.5-78.3) | 75.5 (71.6-78.8) |
4.2 Algorithm Setup
The algorithm setup is now outlined. As discussed earlier, the video modality leverages the D model pre-trained on the action recognition Kinetics dataset [25]. To better suit our ER task, we modified the model’s architecture by adjusting the final linear layer. Specifically, we reconfigured it to output 768-dimensional feature embeddings. This adjustment ensures that both modalities (video and audio) contribute equally-sized feature vectors to the multi-modal representation by fusion through concatenation. The resolution of the RGB video frames was first reduced to pixels. Then, 8 frames from the video stream were randomly selected. To augment the dataset, these frames undergo refinement through random cropping using the TorchVision Library [22], yielding images that enhance the model’s robustness to spatial variations. In addition, we added random horizontal and vertical flips, each with a 30% probability of application, coupled with arbitrary rotations within the range of [-30∘, 30∘].
The audio modality applies an extended version of the HTS-AT model [18], suitable for both multi-channel and single-channel scenarios. As described in Sec. 3, the network structure configuration is arranged into four groups, each containing several Swin-Transformer blocks: 2, 2, 6, and 2, respectively. The mel-spectrogram input is initially transformed into patches and linearly projected to a dimension of . This dimension expands exponentially through each transformer group. It finally reaches a dimension of (), which matches the design principles of Audio Spectrogram Transformer (AST). Pre-processing was carried out as explained in [18] both for multi-channel and single-channel experiments. We augmented the mel-spectrograms by using the SpecAugment Library [21], which consists of temporal masking, occluding four distinct “strips,” each 64 time-frames long. Complementing this, we applied frequency domain masking, obscuring two strips, each 8 frequency bands wide.
Our multi-modal approach combines the feature embeddings from both the video and audio modalities. The 768-dimensional feature vectors extracted from the D model and the extended HTS-AT model are concatenated, resulting in a 1536-dimensional feature representation. This combined feature vector is then fed into a classification head for prediction. The right-hand side of Fig. 1 presents two fully connected layers with a Relu activation function between them, forming the sequence:
The fine-tuning processes are applied using the Adam optimizer with a learning rate of and a warm-up strategy. We used cross-entropy loss as the metric with a batch size of 32. The maximum number of epochs was set to 500 for all experiments, with an early stopping strategy with a patience of 12 to prevent overfitting. In practice, the maximum number of epochs was never reached, as the fine-tuning process was halted earlier due to the activation of the patience parameter. The overall number of parameters for the fine-tuned models are as follows: 32.3M for the uni-modal scheme based on video, 28.7M for the uni-modal scheme based on audio, and 62.7M for the multi-modal scheme.
4.3 Results
Table LABEL:ACE_results details the Accuracy of the various emotion recognition schemes, for seven different rooms from the ACE database. The video-only modality is compared with the audio-only modality (both single- and multi-channel models) and the combined multi-modal approach. As the video modality is unaffected by the acoustic conditions, we only report the results once. To assess the reliability of our results, we report the mean results together with 75% confidence intervals.333github.com/luferrer/ConfidenceIntervals Analyzing Table LABEL:ACE_results, it is observed that for the audio-only schemes, the multi-channel processing methods (Avg mel and Sum PE) consistently outperform the single-channel approaches. This is inline with the findings of [18]. Notably, the multi-modal approaches significantly outperform their uni-modal counterparts. These results are also visually demonstrated in Fig. 2, demonstrating the advantages of the multi-modal processing.
5 Conclusions
In this paper, we presented a MER system designed to operate in reverberant and noisy acoustic environments. Our approach demonstrates robust performance across a range of realistic acoustic conditions by combining an extended multi-channel HTS-AT for audio processing with an D model for video analysis. The MER system combines audio and visual modalities, consistently outperforming uni-modal approaches. By using synthetic RIRs for training and real-world RIRs from the ACE database for testing, we provide a comprehensive assessment of our system’s performance in diverse acoustic environments. Moreover, the utilization of multi-channel audio processing, particularly the Patch-Embed summation, proves beneficial in mitigating the effects of reverberation and noise over the single-channel case. This leads to the potential of our MER system for applications in various real-world scenarios where acoustic conditions may be far from ideal. Future work could focus on further improving the system’s performance in extremely reverberant environments and exploring its effectiveness in other emotional datasets.